PKTOGF(1) PKTOGF(1) NAME pktogf−convertpackedfontfilestogenericfontfiles SYNOPSIS pktogf[−v]pkfilename[gffilename] DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualWeb2C:ATeXimplementation. Thepktogfprogramconvertsapackedfontfile(pk)toagenericfontfile(gf).Packedfontfiles aremuchsmallerthanthecorrespondinggffiles,butsomeDVIreadersonlyunderstandgffiles. Thepkfilenameonthecommandlinemustbecomplete.Becausetheresolutionispartofthe extension,itwouldnotmakesensetoappendadefaultextensionasisdonewithTEXorDVIreadingsoftware.Thegffilenamedefaultstothesame(stripped)nameaspkfilename,anditis placedinthecurrentworkingdirectorywiththegfsuffixreplacingpk.Forexample,theinputfile io.300pkwouldbecomeio.300gf. OPTIONS Unlessthe−vswitchisgiven,pktogfoperatessilently.With−v,theversionnumberoftheprogramandstatisticsabouttheunpackingareoutputtostdout. ENVIRONMENT pktogflooksforpkfilenameusingtheenvironmentvariablePKFONTS.Ifthatisnotset,it usesthevariableTEXPKS.Ifthatisnotset,itusesTEXFONTS.Ifthatisnotset,itusesthe systemdefault. Seetex(1)forthedetailsofthesearching. SEEALSO gftopk(1),gftype(1),pktype(1). DonaldE.Knuthetal.,METAFONTware. AUTHORS TomasRokickiwrotetheprogram.PierreMacKayadapteditforcompilationwithweb2c. Web2C7.32x 14December1993 1